myininaya (myininaya):

I know this is probably not a priority, but I think having more colors to draw with would help us better illustrate somethings to people.

myininaya (myininaya):

Maybe two choices in color like red and black.

OpenStudy (anonymous):

totally agree, and don't forget to add eraser too :P

OpenStudy (karatechopper):

@myin i totally agree @tomas there is an undo button but an eraser would sound good

OpenStudy (asnaseer):

Good idea - the drawing tool really needs a lot of improvement.

OpenStudy (shadowfiend):

Good suggestion. It really does, we put it out and haven't gotten a chance to give it any more love since we initially released it. We're hoping to give it some improvement sometime soon.

OpenStudy (karatechopper):

:)

OpenStudy (mathteacher1729):

http://mugtug.com/sketchpad/ is pretty awesome. Join forces with http://www.twiddla.com !

myininaya (myininaya):

thats really cool mathteacher thanks for the links

myininaya (myininaya):

So this can be seen by more than one person?

myininaya (myininaya):

the sketchpad?

OpenStudy (anonymous):

in twiddia multiple people can draw at the same time

myininaya (myininaya):

twiddla doesn't cost?

myininaya (myininaya):

oh wait it says a free 30 day trail

myininaya (myininaya):

Hey this is off topic, but since I have math teacher here I would like to ask something...

OpenStudy (mathteacher1729):

Go 'right ahead. :)

myininaya (myininaya):

Ok someone wants me to give them a crash course in fractions. I guess the operation of fractions. It is for a job or something. What kinds questions do you think jobs find important about fractions? I know percentages are a big thing.

OpenStudy (anonymous):

yeah i see 30days trial but I think you can use it more than 30 days, how they track it? by cookies?

myininaya (myininaya):

I picture word problems for some reason.

myininaya (myininaya):

I don't know what type of job it is

myininaya (myininaya):

I guess like one of those without a college degree type of jobs

OpenStudy (mathteacher1729):

It depends on the job, really. I mean, off the top of my head, every citizen should know how to: know how to convert basic stuff like 1/2 = .5 and 3/4 = .75 etc know how to add simple fractions 1/2 + 1/3 etc. Instead of teaching the mechanics... "find a common denominator, etc." DRAW A PICTURE and explain what the common denominator MEANS in terms of 1/2 = 5/10 = 50/100 etc. know how to find the percent of a percent of something (3% of the first 20% of $500) know how to have a basic sense of number (THIS is tricky and usually can't be taught in a day or two...) I mean 1/2 + 1/4 , even if you don't know what that is off the top of your head, the sum is LESS THAN ONE. (again, think of a basic picture) When explaining USE MONEY as a metaphor, especially if your friend is old enough to have paid bills before. 2% of $12,870 is ...? What about 2/3 of a 9 hour day on a project? etc. I'd see what they already know about fractions and go from there. Check out http://www.sosmath.com/algebra/fraction/frac1/frac1.html for examples and notes, etc. Good luck to your friend!